Find the unknown side of the right triangle. B . AC 24 вс-? AB 25 |(Round to the nearest thousandth as needed.) Вс-

Find the unknown side of the right triangle. 

AC = 24 

BC = _______ 

AB = 25
Answer #1

we elall tha folroiug tuar Pythagoras bou Ag2+BeAe --0 Now He 25 24 Ae 24 AB 25 tiaugle ABe we have t fnd Be. The ttaaun ABAe
Answer #2

SOLUTION :


The triangle ABC is a right angle one, so,


BC^2 = AB^2 - AC^2

= 25^2 - 24^2

= 49

=> BC = sqrt(49) 

=> BC = 7 (ANSWER).


answered by: Tulsiram Garg